OpenStudy (brooke1556):

Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form. y = x + 8 y = x − 8 y = −5x − 10 y = −5x + 20

OpenStudy (michele_laino):

by definition, the slope \(m\) of the line is: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{10 - 5}}{{2 - \left( { - 3} \right)}} = ...?\]

OpenStudy (brooke1556):

\[\ \frac{ 5 }{ 5 }\]

OpenStudy (welshfella):

yes which simplifies to ?

OpenStudy (brooke1556):

1?

OpenStudy (michele_laino):

correct! next the equation of the line is: \[y - 5 = m\left( {x - \left( { - 3} \right)} \right)\] wherein \(m=1\) please substitute and simplify

OpenStudy (welshfella):

righ so the slope (m) = 1

OpenStudy (brooke1556):

I got 8? @Michele_Laino

OpenStudy (michele_laino):

please write the equation you got

OpenStudy (brooke1556):

\[y-5=1(x-(-3)) \] \[y-5=1x-3\] \[\frac{ 8 }{ 1x } =\frac{ 1x }{ 1x}\] 8

OpenStudy (michele_laino):

we have this: \(y-5=x+3\) next I add \(5\) to both sides, so we get: \(y-5+5=x+3+5\) please simplify

OpenStudy (michele_laino):

please keep in mind that \(x-(-3)=x+3\)

OpenStudy (michele_laino):

hint: what is \(-5+5=...?\) and \(3+5=...?\)

OpenStudy (brooke1556):

0 and 8

OpenStudy (michele_laino):

correct! so the requested equation, is: \(y=x+8\)